# Partition function zeros of quantum many-body systems

### v2

@@ 7 Concluding remarks
For repulsive U, the virtual energies are real in the lowest order calculations. Assuming that a substantial fraction of the virtual roots are real at low $T$ in certain energy ranges, their level density should control the compressibility, heat capacity and susceptibility for the corresponding densities. In this sense the methods developed here may be expected to connect with the Landau Fermi liquid and its many variants[24]. Towards this end it seems important to study the virtual energy distribution for larger systems, and to study its variation with the strength of the interaction $U$, and the possible signatures of a Mott-Hubbard gap. For the case of attractive interaction, the results presented here give non-real virtual energies, already to first order in U and for very small lattice sizes. Further applications of these ideas will be published in a forthcoming publication[25].
@@ Appendix A Zeros from the Luttinger-Ward formula
In this section we discuss another expression for the partition function of a many-body system found in the well-known work of Luttinger and Ward (LW)[20]. LW found a formally exact expression for the thermodynamic potential $\Omega$ (and therefore the partition function) from perturbation theory to all orders in the coupling. The LW result follows from a clever rearrangement of the thermal perturbation theory, whose convergence is implicitly assumed to be sufficiently “benign”. Their result for the grand potential $\Omega$
@@
where $G_{0}$ is the non-interacting Greens function, $\widetilde{\Sigma}^{(\nu)}_{\sigma}(\vec{k},i\omega_{r})$ is the total $\nu$th order skeleton self-energy part, where the order of the diagram $\nu$ counts all explicit powers of the interaction, say U. In these expressions the term $e^{i\omega_{r}0^{+}}$ is required for convergence of the sum. Note that $Z_{B}$ and $Z_{C}$ vanish in the limit of free Fermions. Readers might recognize that $\log Z_{C}$ Eq. (73) is the Luttinger-Ward functional of $G$, with the property that its functional derivative gives the exact self energy, and thereby $\Omega$ is a stationary functional of $G_{\vec{k}{\sigma}}$.

### v3

@@ 7 Concluding remarks
For repulsive U, the virtual energies are real in the lowest order calculations. Assuming that a substantial fraction of the virtual roots are real at low $T$ in certain energy ranges, their level density should control the compressibility, heat capacity and susceptibility for the corresponding densities. In this sense the methods developed here may be expected to connect with the Landau Fermi liquid and its many variants[27]. Towards this end it seems important to study the virtual energy distribution for larger systems, and to study its variation with the strength of the interaction $U$. For the case of attractive interaction, the results presented here give non-real virtual energies, already to first order in U and for very small lattice sizes. Further applications of these ideas will be published in a forthcoming publication[20].
@@ Appendix A Zeros from the Luttinger-Ward formula
In this section we discuss another expression for the partition function of a many-body system found in the well-known work of Luttinger and Ward (LW)[23]. LW found a formally exact expression for the thermodynamic potential $\Omega$ (and therefore the partition function) in terms of the Greens function $G_{\sigma}(\vec{k},i\omega_{r})$ , from perturbation theory to all orders in the coupling. The LW result follows from a clever rearrangement of the thermal perturbation theory, whose convergence is implicitly assumed to be sufficiently “benign”. Their result for the grand potential $\Omega$
@@
where $G_{0}$ is the non-interacting Greens function, $\widetilde{\Sigma}^{(\nu)}_{\sigma}(\vec{k},i\omega_{r})$ is the total $\nu$th order skeleton self-energy part, where the order of the diagram $\nu$ counts all explicit powers of the interaction, in this case U. In these expressions the term $e^{i\omega_{r}0^{+}}$ is required for convergence of the sum. Note that $Z_{B}$ and $Z_{C}$ vanish in the limit of free Fermions. Readers might recognize that $\log Z_{C}$ Eq. (74) is the Luttinger-Ward functional of $G$, with the property that its functional derivative gives the exact self-energy, and thereby $\Omega$ is a stationary functional of $G_{\vec{k}{\sigma}}$.